skip to content
ATNF VLBI Wiki
User Tools
Admin
Log In
Site Tools
Search
Tools
Show pagesource
Old revisions
Export to PDF
Backlinks
Recent Changes
Media Manager
Sitemap
Log In
>
Recent Changes
Media Manager
Sitemap
You are here:
start
»
lbaops
»
actionslba-ops-2016
Trace:
•
actionslba-ops-2016
Sidebar
VLBI Schedule
LBA Home Page
Overview
Documents
Observing
LBA Correlation
Active proposals
eVLBI
Observer's Wiki
Schedules
Sensitivity Calculator
Other VLBI
Edit sidebar
lbaops:actionslba-ops-2016
VLBI/eVLBI operations and technical meeting, Hobart, 15-16 Feb 2016
ACTIONS list
Investigate moving Hobart/Ceduna fully to DBBC (JamieMc)
Aim March 2017
Encourage PIs/Co-Is to come for longer periods (Chris/Phil)
to support other projects in the same block
Find mechanisms to encourage PI observing
we have had some success
Phil says to schedule further in advance
Cormac recommends a direct, strongly worded email
Simon suggests the “withdrawal of future observing time” stick, but we need to say it to all projects, and not have special cases
Sort Tid support issues (Phil)
not being much of an issue, although no host country time support is guaranteed
Tid VNC setup (Chris/Shinj) [done 17/2] (it has come back again)
Tidvsi disk storage (Chris/Shinji) (done)
Check if DBBC can be plugged directly into CA06 (Chris) [done 17/2]
Check if Mopra narrow band filters can be bypassed (Chris) [done 17/2, unknown outcome]
Investigate ARC/Lief wideband proposal (Simon)
Simon has investigated, the proposal has not happened
have to start talking about it November
Clarify standard modes for PIs and document. Review call for proposals(Chris)
not done yet
unclear whether this needs to be done, with most observers using canned setups
Auscope added to lba_monitor (JamieS/Jim)
not done yet
Add Mark5s to recmon (Chris)
not done yet
Trial Aspera (Chris) [email sent out 16/2]
not continuing for many reasons, including that normal data moving is “no effort”
Outstanding 2015 action items
Sched support longterm (Chris)
vex2
Tid DASPro entries (Chris)
not done yet
DASPRo 64
MHz
flipping (Chris)
not done yet
AK07 position (Leonid?)
[Emailed 19-Feb]
cannot happen
RtFC
Sub second DIFX (Chris) [Done 29/2/16]
obs.pl timeout fixes (Chris) [Done 29/2/16]
recdir on flexbuf (Jamie^2)
LBA polarisation - update or advertise memos (Chris)
Automatic e-checklist (JamieS)
EVN tutorial comparison for LBA (Cormac→Chris)
EVN analysis guide actually, but not done yet
Tsys data (Chris)
continued
Continue 1921-293/1934-638 ATCA calibration
not being done, still desirable, possibly use ATCA calibrator database as a first reference
lbaops/actionslba-ops-2016.txt
· Last modified: 2017/02/15 10:20 by
ste616
Page Tools
Show pagesource
Old revisions
Backlinks
Export to PDF
Back to top